Can someone please help me.
Recipe calls for 1/4 teaspoon vanilla. A second recipe calls for 1/2tsp. vanilla and a third recipe calls for 1/8 tsp. vanilla. How much vanilla is needed in total.

Answer by checkley71(8403)

1/4+1/2+1/8
THERE ARE 2 DIFFERENT WAYS TO SOLVE THIS PROBLEM:
CONVERT ALL THESE FRACTIONS TO DECIMAL & ADD.
.5+.25+.125=.875 ANSWER.
OR WE CAN FIND A COMMON DENOMINATOR AND ADD.
1/4=2/8
1/2=4/8
1/8=1/8 NOW WE CAN ADD THEM ALL.
(2+4+1)/8=7/8 ANSWER.
IT SOMETIMES DEPENDS ON THE FORM OF THE ANSWER AS TO WHICH TECHNIQUE YOU USE.
HOWEVER THE ANSWERS ARE THE SAME:
7/8=.875